In 1991, the life expectancy of males in a certain country was 68.5 years. In 1996 it was 71.1 years. Let E represent life expectancy in year t and let t represent the number of years since 1991.
The linear function E(t) that fits the data is
E(t)=?t + ?
Use the function to predict the life expectancy of males in 2009?
E(18)=?
Can someone help me with this one I do not understand. I dont understand the year t and let t represent the number of years. Doesnt t already represent a certain year, then it says it is the number of years since 1991? HELP!!!

Answer by oscargut(2103)

E(t)=at+b , we have to find a and b (real numbers)
In 1991, the life expectancy of males in a certain country was 68.5 years.
t represent the number of years since 1991.
then E(0)=68.5 then b=68.5
then
E(t)=at+68.5
Using now:In 1996 it was 71.1 years
E(5)=a5+68.5=71.1 then 5a=2.6 then a=0.52
then E(t)=0.52t+68.5